If theta is an angle in standard position and P(-3,4) is a point of the terminal side of theta, what is the value of sin theta?
Answer by stanbon(75887)

If theta is an angle in standard position and P(-3,4) is a point of the terminal side of theta, what is the value of sin theta?
------
x = -3 ; y = 4
----
So r = sqrt[9+16] = 5
-----
sin(theta) = y/4 = 4/5
=========================
Cheers,
Stan H.
===============